OpenStudy (anonymous):

solve the system of equations x+2y=2 and x=8-2y

OpenStudy (anonymous):

it has a (1) beside of the first equation I dont know if that means anything

OpenStudy (bobbyleary):

Substitute x = 8-2y into the first equation and solve for x. Then use that x solution and plug that into either equation to solve for y. Let me know if you have any issues.

OpenStudy (radar):

That is kind of tricky. Just try it

OpenStudy (anonymous):

ok i got 8-2y+2y=2 so far

OpenStudy (radar):

that is equivalent to 8=2

OpenStudy (radar):

The equations are contradictive.

OpenStudy (anonymous):

ok so the ordered pair would be (8,2)?

OpenStudy (radar):

Tena32 don't try and trick us! lol

OpenStudy (anonymous):

lol im not this is what it says in my work

OpenStudy (radar):

Look at what the system produces. x+2y=2 x+2y=8 they contrtadict each other, can not be solved.

OpenStudy (anonymous):

it has x+2y=2 (1) x=8-2y (2)

OpenStudy (anonymous):

so there is no solution

OpenStudy (anonymous):

what does the (1) and stuff mean

OpenStudy (radar):

Yes but (2) also says x+2y=8..........add 2y to both sides of the equal sign. You put the (1) and (2) to identify the equations (I guess lol)

OpenStudy (anonymous):

lol i have no clue

OpenStudy (anonymous):

im just checking my answers to see if they are right. some of them I just cant get

OpenStudy (radar):

This is like saying x=8, and x=12 what is really x?

OpenStudy (radar):

Do you have an answer key, just curious what the "approved" answer would be.

OpenStudy (anonymous):

no i dont have it, I wont know if its wrong until i submit it

OpenStudy (radar):

I don't see a solution. Let me think a bit.

OpenStudy (anonymous):

oh its says the solution is blah blah or there is no solution

OpenStudy (anonymous):

those are the two choices

OpenStudy (radar):

go for the no solution

OpenStudy (anonymous):

type an ordered pair or there is no solution

OpenStudy (anonymous):

ok thanks.

OpenStudy (radar):

I am thru thinking. If you plotted each of those linear equations you would get two parallel lines that would never intersect. Usually when you plot linear equations asking for a solution, the lines will intersect at the point of the solution.